OpenStudy (anonymous):

The area of a square is 29 km2. Find the length of a side. Round your answer to one decimal place.

OpenStudy (anonymous):

5.39 km

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

np

OpenStudy (anonymous):

What is the perimeter of the quadrilateral below?

OpenStudy (anonymous):

plz :)

OpenStudy (anonymous):

is it 30 :/

OpenStudy (anonymous):

40

OpenStudy (anonymous):

oh units :/

OpenStudy (anonymous):

40 units i guess lol cuz it doesnt show units on the picture

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

any others lol?

OpenStudy (anonymous):

let me see lol :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

What is the area of the rectangle below?

OpenStudy (anonymous):

84

OpenStudy (anonymous):

thank you one more :P

OpenStudy (anonymous):

What is the formula for the area of a rectangle?

OpenStudy (anonymous):

base*hight

OpenStudy (anonymous):

squared

OpenStudy (anonymous):

lol i forgot to put the choices :P

OpenStudy (anonymous):

A. A = lw B. A = s2 C. A = 2(s + h) D. A = 2(s + l)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

ok yay one more :D

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

What is the perimeter of this polygon?

OpenStudy (anonymous):

73.4 in

OpenStudy (anonymous):

yeeeeeeeah i passed my test thanks to you :D

OpenStudy (anonymous):

lol your welcome :)

OpenStudy (anonymous):

(-__-) i have another test lol

OpenStudy (anonymous):

il help :)

OpenStudy (anonymous):

aww ur the best :)

OpenStudy (anonymous):

What is the area of rhombus DEFG if DF = 8 and EG = 13?

OpenStudy (anonymous):

is it 104 :/

OpenStudy (anonymous):

im pretty sure cuz it doesnt give any other measures

OpenStudy (anonymous):

right it just iges those lengths :/

OpenStudy (anonymous):

it has to be 104

OpenStudy (anonymous):

no it was 52 lol im confused lol :P

OpenStudy (anonymous):

ok one min

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

i dont understand how they got that :/

OpenStudy (anonymous):

yeah im confusded im going to the next one :P lets see .....

OpenStudy (anonymous):

The length of a parallelogram's altitude is also called the _____.

OpenStudy (anonymous):

A. height B. base C. diagonal D. median E. vertex

OpenStudy (anonymous):

a

OpenStudy (anonymous):

yeeah :D

OpenStudy (anonymous):

the other one is the same :P

OpenStudy (anonymous):

look

OpenStudy (anonymous):

What is the area of rhombus ABCD if AC = 11 and BD = 7?

OpenStudy (anonymous):

38.5?

OpenStudy (anonymous):

its right but i did not put sq units -_-

OpenStudy (anonymous):

i got that right? lol

OpenStudy (anonymous):

yeaah lol i did not lol :P

OpenStudy (anonymous):

What is the area of trapezoid KLMN?

OpenStudy (anonymous):

do i multiply

OpenStudy (anonymous):

132 square units?

OpenStudy (anonymous):

nooo lol you have to find out three different areas and add them up

OpenStudy (anonymous):

ok im doing A= 1/2 6 18 +15 :/

OpenStudy (anonymous):

noooo lol ok first you find the area of the square 15*8= 120 then the small triangle .5*8/2=2 then the other triangle 2.5*8/2= 10 then you add all of those together

OpenStudy (anonymous):

so i add 10 and 2 :/ ??? im confused lol

OpenStudy (anonymous):

add 10+ 2+ 120

OpenStudy (anonymous):

so 132 ???

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

it sq units :(

OpenStudy (anonymous):

i get that wrong ur right lol

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

What is the area of trapezoid KLMN?

OpenStudy (anonymous):

300 sq units

OpenStudy (anonymous):

yay thanks :D

OpenStudy (anonymous):

Which formula is similar to the formula for the area of a parallelogram?

OpenStudy (anonymous):

A. The area of a triangle B. The area of a circle C. The area of an ellipse D. The area of a rectangle E. The area of a trapezoid

OpenStudy (anonymous):

What is the area of ABCD?

OpenStudy (anonymous):

lol that pic is the same one as the last question hun

OpenStudy (anonymous):

lolit is :P

OpenStudy (anonymous):

i think its this ona hahahhah sorry :P

OpenStudy (anonymous):

this one is correct lol

OpenStudy (anonymous):

lol give me a min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok im guessin it is 33 sq units

OpenStudy (anonymous):

yeaaaaaah it was lol :D ^.^

OpenStudy (anonymous):

ok good lol

OpenStudy (anonymous):

hahah the other also is the sam e:P

OpenStudy (anonymous):

What is the area of LMNO?

OpenStudy (anonymous):

209 sq units?

OpenStudy (anonymous):

lol 114 is hte answer lol

OpenStudy (anonymous):

how lol?